OpenStudy (anonymous):

a compact disc spins at 2.3 revolutions per second. an ant is walking on the CD and finds that it just begins to slide off the CD when it reaches a point 3.1cm from the CD center

OpenStudy (anonymous):

@roadjester can you help us both on this please

OpenStudy (roadjester):

so what are you looking for?

OpenStudy (roadjester):

Is that the whole question?

OpenStudy (roadjester):

@m806texas @vsandrasue

OpenStudy (anonymous):

did it cut off when it reaches a point of 3.1cm from the CD's center.

OpenStudy (roadjester):

HUH?

OpenStudy (anonymous):

a compact disc spins at 2.3 revolutions per second. an ant is walking on the CD and finds that it just begins to slide off the CD when it reaches a point 3.1cm from the CD's center.

OpenStudy (anonymous):

can u help me find the answer?

OpenStudy (anonymous):

@vsandrasue you have to type the question .....like .... whats the coefficient of friction or whats acceleration

OpenStudy (roadjester):

I need the WHOLE question. EVERYTHING that was given to you.

OpenStudy (anonymous):

oh the question is What is the coefficient of friction between the ant and the CD?

OpenStudy (roadjester):



OpenStudy (roadjester):

so is the movement of the ant perpendicular to the CD?

OpenStudy (anonymous):

it doesnt say. was i used this formula Efx= max=mv2/r to find Us

OpenStudy (roadjester):

Ummm, what does that notation stand for?

OpenStudy (anonymous):

maybe im not understanding which formula to use to find?

OpenStudy (roadjester):

You're on a spinning disk so this problem requires angular movement.

OpenStudy (anonymous):

so there is no formula just an answer?

OpenStudy (anonymous):

A compact disc spins at 2.3 revolutions per second. An ant is walking on the CD and finds that it just begins to slide off the CD when it reaches a point 3.1 cm from the CD's center. (a) What is the coefficient of friction between the ant and the CD? Incorrect: Your answer is incorrect. (b) Is this the coefficient of static friction or kinetic friction? static friction kinetic friction Correct: Your answer is correct.

OpenStudy (anonymous):

i sent you the copy off my homework

OpenStudy (roadjester):

\(\huge \vec F=m\vec a_c=m{v^2\over r}={m(r\omega)^2\over r}={\dfrac {mr^2\omega ^2 } r=mr\omega^2}\)

OpenStudy (anonymous):

it says on the empty box i have to fill in( to enter a number)

OpenStudy (roadjester):

And what is the "Incorrect: Your answer is incorrect." and "Correct: Your answer is correct."?

OpenStudy (anonymous):

do i have to find the omega first?

OpenStudy (anonymous):

there is to questions the second answer is correct. the first question was wrong which was the equation i feel i did wrong.

OpenStudy (roadjester):

omega is given. it is your angular velocity which is currently in revolutions per second You need to convert it to radians per second.

OpenStudy (roadjester):

2\(\pi\) radians=1revolution

OpenStudy (anonymous):

ok i did that and the answer i got was .241

OpenStudy (anonymous):

first answer was 144.513/60 this is my radius per/sec and got the answer of .241

OpenStudy (roadjester):

0.241 rad/s?

OpenStudy (roadjester):

you have 2.3 rev/s. to convert it to rad/s, you multiply by 2pi

OpenStudy (anonymous):

this what i did i put 2.3*2pie divide by 60 seconds.

OpenStudy (roadjester):

why did you divide by 60s?

OpenStudy (anonymous):

i thought i needed to to change to radius/per second. i dont know what i did can you help me find the answer.

OpenStudy (roadjester):

not radius per second; radian per second

OpenStudy (roadjester):

radian is another unit to represent degrees

OpenStudy (roadjester):

for a circle, it is 360 degrees a circle is also 2\(\pi\) radians

OpenStudy (anonymous):

ok radian. can u help me get the right answer please. this what i did (2.3rev/sec)(2pie/1rev)(1min/60sec) which equaled (144.513/60) = .241 was that wrong?

OpenStudy (roadjester):

why do you have a time conversion factor? 2.3 rev/s(2\(\pi\) rad/1rev)=14.45 rad/s

OpenStudy (roadjester):

rad/s is the unit for omega

OpenStudy (anonymous):

so what did i do wrong

OpenStudy (roadjester):

don't divide by 60, keep the units at rad/s and that is your unit for omega

OpenStudy (anonymous):

oh ok so its just 144.513 right

OpenStudy (roadjester):

no, 14.4513

OpenStudy (roadjester):

decimal is in the wrong place

OpenStudy (anonymous):

your right and then use the mvsquared/r to get my Us

OpenStudy (roadjester):

what is Us?

OpenStudy (anonymous):

mue

OpenStudy (anonymous):

u know that weird looking u the mue

OpenStudy (roadjester):

ok use this \(\large\mu\)

OpenStudy (anonymous):

sorry it pronounced mu

OpenStudy (roadjester):

backslash, open parentheses, backslash, mu, backslash, close parentheses

OpenStudy (anonymous):

yes, do i need to solve for that to use the radius

OpenStudy (roadjester):

I'm going to be honest, this problem is whack Not sure if this will work but I'm thinking \(\huge F_k=\mu_kn=\mu_kmg=mr\omega^2\), from there \(\huge \mu_k={\dfrac {r\omega^2}g}\)

OpenStudy (roadjester):

Not sure if that will work since this is three-dimensional and angular, not translational

OpenStudy (roadjester):

I get .6606189

OpenStudy (roadjester):

0.6606189

OpenStudy (anonymous):

ok thats what i got let me put in the answer and see if thats correct

OpenStudy (anonymous):

that was correct thank you.

OpenStudy (roadjester):

shouldn't thank me thank @m806texas I think she said you two were classmates.